OpenStudy (faithspring):

The midpoint of UV is (5,-11). The coordinates of one endpoint are U(3,5). Find the coordinates of endpoint V. (please show work and explain.)

OpenStudy (will.h):

\[\frac{ x1 + x2 }{ 2 } , \frac{ y1 + y2}{ 2 } = (5,-11)\] \[\frac{ 3 + x2 }{ 2 } , \frac{ 5 + y2}{ 2 } = (5,-11)\] Admit it This looks beautiful i'll be right back

OpenStudy (faithspring):

what's the answer? @Will.H

OpenStudy (triciaal):

read the response and continue. If you cannot continue what question do you have?

OpenStudy (faithspring):

What do I do next? @triciaal

OpenStudy (triciaal):

the midpoint is 1/2 way between the 2 points do the x separate from the y

OpenStudy (will.h):

That''s also a good way yes indeed midpoint is 1/2 way between the 2 points you may multiply the x and y values of the midpoint by 1/2 and you'll get your other point

OpenStudy (triciaal):